# A probabilistic proof of Euler’s pentagonal number theorem

### v2

@@ 4. Probabilities
While we have determined the event spaces for our investigation, the next task is to analyze the probabilities. Throughout, for a probability model, we denote by $\mathbf{Pr}_{E}$ the total probability of all outcomes in the event space $E$.
Recalling again the definition of the model k, we have
@@
The following evaluation is then immediate.
@@
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\prod_{i\in I}q^{i}\prod_{j\in J}(1-q^{j}).$ (4.1)
@@
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=0.$ (4.2)
@@
Let $I$ and $I^{\prime}$ be two finite subsets of $\mathbb{N}_{+}$, not necessarily disjoint, and let $J$ be a subset of $\mathbb{N}_{+}$ that is disjoint with both $I$ and $I^{\prime}$. Suppose that
@@
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{Pr}_{\langle I^{\prime}_{\triangleright},J_{\triangleleft}\rangle}.$ (4.3)
Finally, we recall that in a probability model, if two event spaces $E_{1}$ and $E_{2}$ are disjoint, then
$\mathbf{Pr}_{E_{1}\sqcup E_{2}}=\mathbf{Pr}_{E_{1}}+\mathbf{Pr}_{E_{2}}.$
@@
Let $I$, $J$ and $K$ be pairwise disjoint subsets of $\mathbb{N}_{+}$. Then
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\sum_{K^{\prime}\subset K}\mathbf{Pr}_{\langle(I\sqcup K^{\prime})_{\triangleright},(J\sqcup(K\backslash K^{\prime}))_{\triangleleft}\rangle}.$ (4.4)
@@
$\mathbf{Pr}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{Pr}_{\langle(I\sqcup\{k\})_{\triangleright},J_{\triangleleft}\rangle}+\mathbf{Pr}_{\langle I_{\triangleright},(J\sqcup\{k\})_{\triangleleft}\rangle}.$ (4.5)
@@ 5. The pentagonal number theorem
The key to our proof of the pentagonal number theorem is the following relation.
@@
$\mathbf{Pr}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}$ $=\sum_{n=0}^{N}(-1)^{n}\big{(}\mathbf{Pr}_{\langle[n+1,2n]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle[n+1,2n+1]_{\triangleright},\varnothing_{\triangleleft}\rangle}\big{)}$
$\quad-(-1)^{N}\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}.$ (5.1)
@@
$\mathbf{Pr}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle\varnothing_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle\{1\}_{\triangleright},\varnothing_{\triangleleft}\rangle}-\sum_{k\geq 1}\mathbf{Pr}_{\langle\{k+1\}_{\triangleright},[1,k]_{\triangleleft}\rangle}$
$=\mathbf{Pr}_{\langle\varnothing_{\triangleright},\varnothing_{\triangleleft}\rangle}-\sum_{k\geq 1}\mathbf{Pr}_{\langle\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\rangle}.$
Note that $\big{\langle}\varnothing_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}$ contains all shuffles of $\mathbb{N}$. Meanwhile, given an arbitrary shuffle, either none of the rounds are leftmost (which is characterized by $\big{\langle}\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\big{\rangle}$) or there exists a unique round $k\geq 1$ such that k is the first leftmost shuffling (which is characterized by $\big{\langle}\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\big{\rangle}$). Furthermore, given two different positive integers $k$ and $k^{\prime}$, we assume without loss of generality that $k<k^{\prime}$. Then $k\in\{k\}$ and $k\in[1,k^{\prime}-1]$. In light of Proposition 3.1,
@@
Applying Proposition 4.4 leads us to the required relation. ∎
@@
It suffices to show that for each nonnegative integer $N$,
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+3+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}.$
We start by noting that for each $k\geq 1$, the sets $[k+1+N,k+1+2N]$, $[2+N,k+N]$ and $\{1+N\}$ are pairwise disjoint. It follows from Proposition 4.4 that
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle(\{1+N\}\sqcup[k+1+N,k+1+2N])_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
For the first sum on the right-hand side,
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad+\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}.$
@@
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}$ $=\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
@@
$\sum_{k\geq 1}\mathbf{Pr}_{\langle(\{1+N\}\sqcup[k+1+N,k+1+2N])_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}=\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
@@
$\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{Pr}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{Pr}_{\langle[k+1+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
Singling out the $k=1$ term of the sum in the above gives the desired part $\mathbf{Pr}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$. ∎
@@
$\mathbf{Pr}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}=\sum_{n\geq 0}(-1)^{n}\big{(}\mathbf{Pr}_{\langle[n+1,2n]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{Pr}_{\langle[n+1,2n+1]_{\triangleright},\varnothing_{\triangleleft}\rangle}\big{)}.$
Applying Proposition 4.1 yields
@@ 6. Epilogue
Before putting down the curtain, let us flash back to Euler’s original proof of the pentagonal number theorem in [3]. Adopting the $q$-Pochhammer symbol for $n\in\mathbb{N}\cup\{\infty\}$:
@@
the key observation of Euler, although not explicitly stated, is as follows:
@@
$\mathbf{Pr}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\prod_{i=k+1+N}^{k+1+2N}q^{i}\prod_{j=1+N}^{k+N}(1-q^{j})$
@@
In view of (5.1), this is clearly the last piece of expressions required for (6.1).
@@ Acknowledgements
This work was supported by the Austrian Science Fund (No. 10.55776/F1002). I am grateful to Hugh Thomas for the comments on Proposition 4.4 in an earlier draft.

### v3

@@ 4. Probabilities
In the previous section, we have determined the desired event spaces $\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}$ attached to the shuffling model $\Sh=(\Omega,\mathbf{P})$ where $I,J\subset\mathbb{N}_{+}$ are disjoint. Now our task is to analyze the probability $\mathbf{P}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}$.
@@
[…]
In particular, this probability evaluation is independent of $\hat{\mathfrak{s}}_{k}$, or equivalently, the outcome $\mathfrak{s}_{k-1}$ at Time k1k-1.
This independence, together with the nature of a Markov process, immediately gives the following probability evaluation for the event space $\big{\langle}I_{\triangleright},J_{\triangleleft}\big{\rangle}$ attached to the shuffling model .
@@
$\mathbf{P}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\prod_{i\in I}q^{i}\prod_{j\in J}(1-q^{j}).$ (4.1)
@@
$\mathbf{P}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=0.$ (4.2)
@@
Let $I$ and $I^{\prime}$ be two finite subsets of $\mathbb{N}_{+}$, not necessarily disjoint, and let $J$ be a subset of $\mathbb{N}_{+}$ that is disjoint from both $I$ and $I^{\prime}$. Suppose that
@@
$\mathbf{P}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{P}_{\langle I^{\prime}_{\triangleright},J_{\triangleleft}\rangle}.$ (4.3)
Finally, we recall that, in a probability model, if two event spaces $E_{1}$ and $E_{2}$ are disjoint, then
$\mathbf{P}_{E_{1}\sqcup E_{2}}=\mathbf{P}_{E_{1}}+\mathbf{P}_{E_{2}}.$
@@
Let $I$ and $J$ be disjoint subsets of $\mathbb{N}_{+}$. Then for any subset $K$ of $\mathbb{N}_{+}$ that is disjoint from both $I$ and $J$,
$\mathbf{P}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\sum_{K^{\prime}\subset K}\mathbf{P}_{\langle(I\sqcup K^{\prime})_{\triangleright},(J\sqcup(K\backslash K^{\prime}))_{\triangleleft}\rangle}.$ (4.4)
@@
$\mathbf{P}_{\langle I_{\triangleright},J_{\triangleleft}\rangle}=\mathbf{P}_{\langle(I\sqcup\{k\})_{\triangleright},J_{\triangleleft}\rangle}+\mathbf{P}_{\langle I_{\triangleright},(J\sqcup\{k\})_{\triangleleft}\rangle}.$ (4.5)
@@ 5. The pentagonal number theorem
The key to our probabilistic proof of the pentagonal number theorem is the following relation.
@@
$\mathbf{P}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}$ $=\sum_{n=0}^{N}(-1)^{n}\big{(}\mathbf{P}_{\langle[n+1,2n]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{P}_{\langle[n+1,2n+1]_{\triangleright},\varnothing_{\triangleleft}\rangle}\big{)}$
$\qquad\qquad\qquad\qquad-(-1)^{N}\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}.$ (5.1)
@@
$\mathbf{P}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}$ $=\mathbf{P}_{\langle\varnothing_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{P}_{\langle\{1\}_{\triangleright},\varnothing_{\triangleleft}\rangle}-\sum_{k\geq 1}\mathbf{P}_{\langle\{k+1\}_{\triangleright},[1,k]_{\triangleleft}\rangle}$
$=\mathbf{P}_{\langle\varnothing_{\triangleright},\varnothing_{\triangleleft}\rangle}-\sum_{k\geq 1}\mathbf{P}_{\langle\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\rangle}.$
Note that $\big{\langle}\varnothing_{\triangleright},\varnothing$$_{\triangleleft}\big{\rangle}$ contains all shuffles of $\mathbb{N}$ so $\big{\langle}\varnothing_{\triangleright},\varnothing_{\triangleleft}\big{\rangle}=\Omega$. Meanwhile, given an arbitrary shuffle $\mathfrak{s}\in\Omega$, we project it onto each model k to get an outcome shuffle $\mathfrak{s}_{k}\in\Omega_{k}$ at Time $k$. Then either every $\mathfrak{s}_{k}$ is produced by an anti-leftmost shuffling (this case corresponds to $\big{\langle}\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\big{\rangle}$) or there exists a unique time $k$ such that $\mathfrak{s}_{k}$ is the first shuffle produced by a leftmost shuffling (this case corresponds to $\big{\langle}\{k\}_{\triangleright},[1,k-1]_{\triangleleft}\big{\rangle}$). Furthermore, given two different positive integers $k$ and $k^{\prime}$, we assume without loss of generality that $k<k^{\prime}$. Then $k\in\{k\}$ and $k\in[1,k^{\prime}-1]$. In light of Proposition 3.4,
@@
Application of Proposition 4.4 leads us to the required relation. ∎
@@
It suffices to show that, for each nonnegative integer $N$,
$\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{P}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{P}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{P}_{\langle[k+2+N,k+3+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}.$
We start by noting that, for each $k\geq 1$, the sets $[k+1+N,k+1+2N]$, $[2+N,k+N]$ and $\{1+N\}$ are pairwise disjoint. It follows from Proposition 4.4 that
$\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{P}_{\langle(\{1+N\}\sqcup[k+1+N,k+1+2N])_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
We split the first sum on the right-hand side as
$\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{P}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad+\sum_{k\geq 1}\mathbf{P}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}.$
@@
$\sum_{k\geq 1}\mathbf{P}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+1+N]_{\triangleleft}\rangle}$ $=\sum_{k\geq 1}\mathbf{P}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
@@
$\sum_{k\geq 1}\mathbf{P}_{\langle(\{1+N\}\sqcup[k+1+N,k+1+2N])_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}=\sum_{k\geq 1}\mathbf{P}_{\langle[k+2+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
@@
$\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\mathbf{P}_{\langle[2+N,2+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$
$\quad-\sum_{k\geq 1}\mathbf{P}_{\langle[k+1+N,k+2+2N]_{\triangleright},[2+N,k+N]_{\triangleleft}\rangle}.$
By singling out the $k=1$ term of the sum in the above expression, we obtain the desired part $\mathbf{P}_{\langle[2+N,3+2N]_{\triangleright},\varnothing_{\triangleleft}\rangle}$. ∎
@@
$\mathbf{P}_{\langle\varnothing_{\triangleright},[1,\infty)_{\triangleleft}\rangle}=\sum_{n\geq 0}(-1)^{n}\big{(}\mathbf{P}_{\langle[n+1,2n]_{\triangleright},\varnothing_{\triangleleft}\rangle}-\mathbf{P}_{\langle[n+1,2n+1]_{\triangleright},\varnothing_{\triangleleft}\rangle}\big{)}.$
Application of Proposition 4.1 yields
@@ 6. Epilogue
Before putting down the curtain, let us flash back to Euler’s original proof of the pentagonal number theorem in [3]. Adopting the $q$-Pochhammer symbol for $n\in\mathbb{N}\cup\{\infty\}$,
@@
the key observation of Euler, although not explicitly stated, is the following.
@@
$\mathbf{P}_{\langle[k+1+N,k+1+2N]_{\triangleright},[1+N,k+N]_{\triangleleft}\rangle}$ $=\prod_{i=k+1+N}^{k+1+2N}q^{i}\prod_{j=1+N}^{k+N}(1-q^{j})$
@@
In view of (5.1), this is clearly the last piece of expression required for (6.1).
@@ Acknowledgements
This work was supported by the Austrian Science Fund (No. 10.55776/F1002). I would like to thank the anonymous referee for the inspiring suggestions that have improved the quality of this paper to a great extent. I am also grateful to Hugh Thomas for the comments on Proposition 4.4 in an earlier draft.